OpenStudy (anonymous):

Find the general solutions of 1 0 -5 0 -8 3 0 1 4 -1 0 6 0 0 0 0 1 0 0 0 0 0 0 0 Last question I'm going to ask, I promise.

OpenStudy (anonymous):

Is there even a solution? The third row doesn't make a lot of sense 0 + 0 + 0 + 0 + 1 = 0?

ganeshie8 (ganeshie8):

is that the augmented matrix in row echelon form ?

OpenStudy (anonymous):

I'm not sure. The problem doesn't say.

OpenStudy (anonymous):

Wait! okay, it is the augmented matrices.

ganeshie8 (ganeshie8):

good, how many pivots are there ?

OpenStudy (anonymous):

I don't know. How do I tell if they are pivots?

ganeshie8 (ganeshie8):

in echelon form, pivot is the first "non zero" number in a row

OpenStudy (anonymous):

So there would be three pivots then.

ganeshie8 (ganeshie8):

Right, `1` 0 -5 0 -8 3 0 `1` 4 -1 0 6 0 0 0 0 `1` 0 0 0 0 0 0 0 you have 3 pivot columns and 2 free columns lets find a particular solution by letting free variables = 0

OpenStudy (anonymous):

Okay.

ganeshie8 (ganeshie8):

just for definiteness, call the variables : x1, x2, x3, x4, x5 so to find a particular solution you let the free variables x3 = x4 = 0 and solve for the pivot variables x1, x2, x5

ganeshie8 (ganeshie8):

x1 x2 x3 x4 x5 b `1` 0 -5 0 -8 3 0 `1` 4 -1 0 6 0 0 0 0 `1` 0 0 0 0 0 0 0 ^ ^ free columns

ganeshie8 (ganeshie8):

notice that the corresponding system is ``` x5 = 0 x2 + 4x3 - x4 = 6 x1 - 5x3 - 8x5 = 3 ``` let x3 = x4 = 0 and solve remaining variables

OpenStudy (anonymous):

Okay, so how do I solve for the pivot points?

ganeshie8 (ganeshie8):

plugin x3=x4=0 in above system

ganeshie8 (ganeshie8):

then you will have 3 equations and 3 unknowns in triangular form which is a piece of cake to solve by back substitution

OpenStudy (anonymous):

so.. x2 = 5 x5 = 0 x1=0

ganeshie8 (ganeshie8):

try again

ganeshie8 (ganeshie8):

the corresponding system is ``` x5 = 0 x2 + 4x3 - x4 = 6 x1 - 5x3 - 8x5 = 3 ``` letting x3 = x4 = 0, the system becomes ``` x5 = 0 x2 = 6 x1 - 8x5 = 3 ``` yes ?

OpenStudy (anonymous):

Yes.

ganeshie8 (ganeshie8):

solve it

ganeshie8 (ganeshie8):

its trivial, but you have to do it

OpenStudy (anonymous):

I'm trying to do it right now, give me a moment.

OpenStudy (anonymous):

x5 = 0 x2 + 4(0) - 0 = 6 -> x2 = 6 x1 - 5(0) - 8(0) = 3 -> x1 = 3

ganeshie8 (ganeshie8):

so \(\begin{pmatrix} 3\\6\\0\\0\\0 \end{pmatrix}\) is a particular solution we still need to find the null solution, familiar with the process ?

OpenStudy (anonymous):

Nope

ganeshie8 (ganeshie8):

null solution is the solution to following system : ``` x5 = 0 x2 + 4x3 - x4 = 0 x1 - 5x3 - 8x5 = 0 ```

ganeshie8 (ganeshie8):

Notice that I have put 0s on the right hand side now

ganeshie8 (ganeshie8):

again we mess with the free variables to cookup the null solution : since we have two free variables, we expect to get two independent null solutions

OpenStudy (anonymous):

Do I use the values for the system that we came up with earlier?

ganeshie8 (ganeshie8):

let x3 = 0, x4 = 1 and find one null solution let x3 = 1, x4 = 0 and find the other null solution

ganeshie8 (ganeshie8):

null solution is the solution to following system : ``` x5 = 0 x2 + 4x3 - x4 = 0 x1 - 5x3 - 8x5 = 0 ``` plugin x3 = 0, x4 = 1 and solve the other variables

OpenStudy (anonymous):

x2 - 1 = 0 x2 = 1

ganeshie8 (ganeshie8):

yes, keep going find the remaining two variables

OpenStudy (anonymous):

x1 = 0

OpenStudy (anonymous):

Unless you wanted me to use "let x3 = 1, x4 = 0 and find the other null solution" to solve x1 - 5x3 - 8x5 = 0

OpenStudy (anonymous):

Because then x1 would equal 5

ganeshie8 (ganeshie8):

one thing at a time

ganeshie8 (ganeshie8):

null solution is the solution to following system : ``` x5 = 0 x2 + 4x3 - x4 = 0 x1 - 5x3 - 8x5 = 0 ``` plugin x3 = 0, x4 = 1 and solve the other variables this gives a null solution \(\begin{pmatrix} 0\\1\\0\\1\\0\end{pmatrix}\), yes ?

OpenStudy (anonymous):

Yeah.

ganeshie8 (ganeshie8):

save it, find the other null solution also similarly

ganeshie8 (ganeshie8):

null solution is the solution to following system : ``` x5 = 0 x2 + 4x3 - x4 = 0 x1 - 5x3 - 8x5 = 0 ``` plugin x3 = 1, x4 = 0 and solve the other variables

OpenStudy (anonymous):

Okay, so then x1 = 1 and x2 = -1

ganeshie8 (ganeshie8):

try again

OpenStudy (anonymous):

Thanks for your help ganeshie8, but I have to go to sleep, it's really late here and I have to wake up for class. Have a wonderful evening, and I'm sorry for not being about to finish the problem.

ganeshie8 (ganeshie8):

we're mostly done just wait one minute

ganeshie8 (ganeshie8):

null solution is the solution to following system : ``` x5 = 0 x2 + 4x3 - x4 = 0 x1 - 5x3 - 8x5 = 0 ``` plugin x3 = 1, x4 = 0 and solve the other variables that gives the null solution \(\begin{pmatrix} 5\\-4\\1\\0\\0\end{pmatrix}\)

ganeshie8 (ganeshie8):

so the nullspace is given by \[\large c_1\begin{pmatrix} 0\\1\\0\\1\\0\end{pmatrix} + c_2\begin{pmatrix} 5\\-4\\1\\0\\0\end{pmatrix}\]

ganeshie8 (ganeshie8):

then the final general solutions is given by \(x_p + x_{null}\) : \[x_{\text{general }} = \begin{pmatrix} 3\\6\\0\\0\\0 \end{pmatrix} +\large c_1\begin{pmatrix} 0\\1\\0\\1\\0\end{pmatrix} + c_2\begin{pmatrix} 5\\-4\\1\\0\\0\end{pmatrix} \]

OpenStudy (anonymous):

Closing now as it is no longer helpful.